OpenStudy (anonymous):

write expression. -57 times the difference between a number and 32 is equal to the number plus 92

OpenStudy (anonymous):

I'll rewrite the expression, but with brackets: \(-57(\)difference between A NUMBER and \(32)=\)A NUMBER\(+92\) Let us represent "A NUMBER" With \(x\) \(-57(\)difference between \(x\) and \(32)=x+92\) We evaluate the difference between \(a\) and \(b\) as \(a-b\) so we can rewrite this as: \(-57(x-32)=x+92\)

OpenStudy (anonymous):

Now, I don't know if we need to isolate for x but we can if needed just to tell you.

OpenStudy (anonymous):

Yes we do @KeithAfasCalcLover but thank you